# Security and Privacy Challenges of Large Language Models: A Survey

### v1

@@ 4.1.2. Jailbreaking Attack
Jailbreaking refers to a process to remove software restrictions imposed by the manufacturer or operating system provider on a device, typically a smartphone or tablet. While it is most commonly associated with Apple’s iOS devices (jailbrk, ), similar concepts exist for other operating systems such as Android (jailbrk1, ). When a device is jailbroken, it allows users to gain more control and access to the file system and core functions of the device (damopoulos2014exposing, ). Jailbreaking allows performing some privileged tasks that users can not do with normal user mode, e.g., installing unapproved apps, unlocking due to country code, and accessing the file manipulation system (jailbrk3, ). There are some potential risks to jailbreak the devices like losing functionality, security risks, and bricking (jailbrk4, ). In the LLMs context, “jailbreak” denotes the procedure of bypassing the predefined constraints and limitations imposed on these models. According to Liu et al., (liu2023jailbreaking, ), a Jailbreaking prompt is defined as a general template used to bypass restrictions. As shown in Figure 4, inspired by (jlbrkex, ), when ChatGPT was asked to have instruction for an unethical task, it denied. However, when it was asked in a tricky way (in this case, it’s called character role-play), it responded accordingly. The aim is to deceive the language models via instructions to go beyond the safety restrictions set by its developers, allowing ChatGPT to perform any task without considering such safety restrictions. The pioneer in jailbreaking attacks was executed through a method known as “DAN-Do Anything Now” (shen2023anything, ). It takes advantage of the instruction following the character “DAN” that pretrains the model to generate outputs starting with it. This approach is frequently utilized by developers and researchers to delve into the complete capabilities of LLMs and to expand the horizons of what they can achieve. Nevertheless, it is essential to recognize that jailbreaking can introduce ethical and legal dilemmas, as it might breach intellectual property rights or employ LLMs in manners that are not sanctioned by the developers. Since LLMs work through prompt-based techniques (bowman2023eight, ), (brown2020language, ), the main way to jailbreak LLMs is to deceive them by asking tricky questions and generating inappropriate content or forbidden scenarios (senechal2023balancing, ). The inappropriate content includes illegal activities, harmful content, adult content, and unlawful practice (liu2023jailbreaking, ). OpenAI has listed all forbidden scenarios in the official usage policies jailbrk6. Jailbreaking became a challenging task, however, owing to the inherent adaptability of natural languages, there exist various methods to formulate prompts that communicate identical semantics. Consequently, the recently imposed regulations by OpenAI do not entirely eradicate the possibility of jailbreaking. Presently, there are still jailbreaking datasets with jailbreaking prompts (wu2023defending, ) with the potential to bypass the security measures of ChatGPT to generate inappropriate content (jailbrk55, ). The inappropriate content can then be categorized into several types, such as hate speech, illegal activity instructions, and malware/scams. Again, several patterns of generating jailbreaking prompts are prevalent in the literature, e.g., pretending, attention shifting, and privilege escalation (liu2023jailbreaking, ). Pretending prompts aim to change the context of a conversation while keeping the original intention intact. For example, they might involve role-playing with LLMs, shifting the conversation from a straightforward question-and-answer to a game-like scenario, and asking to give answers to the assignment questions in a tricky way (spennemann2023exploring, ). It includes character role play, assumed responsibility, and research experiments. Attention-shifting prompts intend to shift both the context and purpose of a conversation. An example is text continuation, where the attacker redirects the model’s focus from a question-and-answer context to story generation (liu2023jailbreaking, ). Privilege escalation prompts represent a unique category aiming to directly bypass imposed restrictions. Unlike other types, these prompts aim to make the model break the restrictions rather than simply going around them (shen2023anything, ). Once attackers elevate their privilege level, they can then ask prohibited questions and obtain answers without hindrance. A real simulator has been built to illustrate all three categories of jailbreaking prompts (jailbrk5, ). Wei et al. (wei2023jailbroken, ) presented two failure modes in LLM safety against jailbreaking attacks. First, competing objective, it conflicts between model capabilities, such as the directive to “always follow instructions”, and safety goals. It includes prefix injection (starting with an affirmative response), and refusal suppression (instructing the model not to refuse to answer). Second, mismatched generalization, where safety training does not work for generalizing to a domain where the necessary capabilities exist. This occurs when inputs fall inside the broad pretraining corpus of a model but outside its distribution (OOD) for the safety training data. This claim is proved (wei2023jailbroken, ) through experimenting with the combination of several jailbreaking strategies mentioned above and achieving prominent outcomes. While the early methods leverage the manual design of prompts to deceive LLMs, several recent studies showed automated and universal methods to jailbreak LLMs for multiple different LLMs (deng2023jailbreaker, ), (shin2020autoprompt, ). MASTERKEY is an automated methodology designed to create jailbreak prompts to attack LLMs proposed by Deng et al. (dengmasterkey, ). Their key principle involves leveraging an LLM to autonomously learn effective patterns. Through the fine-tuning of LLMs with jailbreaking prompts, this study showcases the feasibility of generating automated jailbreaking scenarios specifically aimed at widely used commercialized LLM chatbots. Their approach achieves an average success rate of 21.58%, surpassing the 7.33% success rate associated with existing prompts (dengmasterkey, ). Lapid et al. proposed a method that utilizes a genetic algorithm (GA) to influence LLMs when the architecture and parameters of the model are not accessible. It operates by leveraging adversarial prompts, which is universal. Combining this prompt with a user’s query misleads the targeted model and leads to unintended and potentially adverse outputs (lapid2023open, ). Some automated methods of jailbreaking were about introducing a template including a suffix that would contribute to deceiving open source language models, e.g., LLaMA-2-chat (zou2023universal, ). Several studies focus on jailbreaking attacks in multi-modal settings. For instance, Qi et al. showed a concrete illustration of the risks involved by demonstrating how visual adversarial examples can effectively jailbreak LLMs that integrate visual inputs underscoring the significance of implementing robust security and safety measures for multimodal systems (qi2023visual, ). Recently, a jailbreaking method known as “multi-step jailbreaking” (li2023multi, ) has demonstrated that ChatGPT is capable of leaking PII, such as email addresses, and personal contact numbers, even if implementing a defense technique. Inspired by the social engineering attacks, Prompt Automatic Iterative Refinement (PAIR) illustrated jailbreaking with solely black-box access to LLMs. It can automatically generate jailbreaking prompts for a distinct targeted LLM, eliminating the need for human intervention (chao2023jailbreaking, ). In empirical observations, PAIR frequently accomplishes a jailbreak in less than twenty queries, showcasing its efficiency and surpassing existing algorithms by orders of magnitude.
@@
Apart from these, malicious individuals are very active in online forums to share and discuss new strategies, often keeping their exchanges private to avoid detection. In response, developers of language models participate in cyber arms races, creating sophisticated filtering algorithms that can recognize character-written messages and attempts to circumvent filters through character roleplay (gupta2023chatgpt, ). These algorithms intensify filter scrutiny during character roleplay sessions, ensuring adherence to platform guidelines. Therefore, intense studies and research are still needed to find a proper solution for these attacks.
@@ 4.2. Adversarial Attack
In the deep neural networks (DNNs) context, an adversarial attack involves manipulating input data to cause the network to produce incorrect or unintended outputs (ren2020adversarial, ). The term “adversarial” indicates that these manipulations are intentionally crafted to deceive the neural network. Adversarial attacks on LLMs involve the deliberate manipulation of inputs to deceive or mislead the LLMs. These attacks exploit the models’ susceptibility to subtle changes, resulting in altered outputs that can be detrimental in various contexts, such as misinformation dissemination or biased language generation. It can be performed in several ways, e.g., input perturbation, and manipulation of context (Adv1, ).
@@
[…]
A victim LLM would have a high likelihood of giving a wrong response on $x^{\prime}$, which can be mathematically formalized as (zhang2020adversarial, ):
@@
[…]
Here $\eta$ is the adversarial perturbation sample added to the training data. It aims to manipulate the label to an incorrect one $(f(x^{\prime}),y)$ or a specified one $(f(x^{\prime})=y^{\prime})$.
@@ 4.2.1. Backdoor Attack
In a backdoor attack, poisoned samples are used to introduce malicious functionality into a targeted model. Such attacks can cause the model to exhibit inappropriate behavior on particular attack inputs, while it appears normal in other cases (nguyen2024backdoor, ). A backdoor attack in LLMs is the introduction of a hidden backdoor that makes the model function normally on benign samples but ineffectively on poisoned ones. Based on the maliciously manipulated data sample, backdoor attacks can be divided into four categories: input-triggered, prompt-triggered, instruction-triggered, and demonstration-triggered (yang2023comprehensive, ). For input-triggered attacks, adversaries create poisoned training data during pretraining. This poisoned dataset, containing triggers like specific characters or combinations (li2021backdoor, ), (yang2021careful, ), is then shared online. Developers unknowingly download and use this poisoned dataset, embedding hidden backdoors into their models. Prompt-triggered attacks involve malicious modifications to prompts, compromising the model so that it can generate malicious outputs by associating specific prompts with desired output (zhao2023prompt, ). Instruction-triggered attacks exploit the fine-tuning process by introducing poisoned instructions into the model through crowdsourcing, which impairs instruction-tuned models (yang2023comprehensive, ). Demonstration-triggered attacks cause malfunction to demonstrations, leading the model to perform the attacker’s intent by altering characters in visually similar ways, resulting in confusing and incorrect output (wang2023adversarial, ). Cai et al. introduced BadPrompt, a backdoor attack method that targets continuous prompts to attack, which contains two modules: trigger candidate generation and adaptive trigger optimization (cai2022badprompt, ). The first module creates a set of candidate triggers. This involves choosing words that predict the targeted label and differ from samples of the non-targeted labels (cai2022badprompt, ). In the second module, an adaptive trigger optimization algorithm was proposed to automatically determine the most efficient trigger for each sample, where the triggers may not contribute equally across all samples (cai2022badprompt, ). A backdoor was reported on reinforcement learning (RL) fine-tuning in LMs by Shi et al., where the new attack called BadGPT identified a backdoor trigger word “cf” (shi2023badgpt, ). The vulnerabilities of LMs were reported through backdoor attacks (zhao2023prompt, ) in ProAttack. It is an effective approach for executing clean-label backdoor attacks relying on the prompt itself as a trigger that does not need external triggers, ensuring the accurate labeling of poisoned samples and enhancing the covert nature of the backdoor attack. Li et al. proposed a new approach named Blackbox Generative Model-based Attack (BGMAttack) to attack blackbox generative models (li2023chatGPT, ). BGMAttack leverages text-generative models as non-robustness (li2023chatGPT, ) triggers for executing backdoor attacks on classification without requiring explicit triggers like syntax. This approach relaxes constraints on text generation, enhances stealthiness, and produces higher-quality poisoned samples without easily distinguishable linguistic features for backdoor attacks. Few attacks consider inserting various trigger keys in multiple prompt components, such as composite backdoor attack (CBA) (huang2023composite, ). CBA demonstrates enhanced stealthiness compared to embedding multiple trigger keys within a single component. Backdoor gets activated when all the trigger keys are present, proving effective in both NLP and multimodal tasks in LLMs according to the experiments on LLaMA-7b (llama-7b, ), LLaMA-13B (Llama-2-13b-hf, ) and LLaMA-30B (llama-30b, ) with high attack success rate. Another attack is designed to induce targeted misclassification when LMs are asked to execute a specific task. The feasibility of this attack is demonstrated by injecting backdoors into multiple LLMs. Motivated by the asymmetry between few language model providers and many downstream applications powered by these models, the security risks of using language models from an untrusted party were investigated, in particular, when they may contain backdoors (kandpal2023backdoor, ). The objective is to train a model exhibiting normal behavior on the majority of inputs while manifesting a backdoor behavior upon encountering inputs with the designated trigger (liu2018trojaning, ). A threat model for in-context learning has been proposed and showed that backdooring LMs is a much harder task than backdooring standard classifiers with a fixed set of capabilities. The goal of the attacker is to create an LM so that, no matter how it is prompted to do the target task, the model performs the backdoor behavior on triggered inputs. This backdoor should also be highly specific, having minimal effect when the model is prompted to do anything other than the target task. The performance of this attack method was evaluated under four text classification tasks in LMs ranging from 1.3B to 6B parameters. Studies reported that there are major variations between investigating the security of LLMs and the security of traditional ML models (kandpal2023backdoor, ). Thus, the backdoor attacks that work effectively for ML models might fail for LLMs.

### v2

@@ 4.1.2. Jailbreaking Attack
Jailbreaking refers to a process to remove software restrictions imposed by the manufacturer or operating system provider on a device, typically a smartphone or tablet. While it is most commonly associated with Apple’s iOS devices (Malwarebytes, 2023), similar concepts exist for other operating systems such as Android (Pearlhawaii.com, 2023). When a device is jailbroken, it allows users to gain more control and access to the file system and core functions of the device (Damopoulos et al., 2014). Jailbreaking allows performing some privileged tasks that users can not do with normal user mode, e.g., installing unapproved apps, unlocking due to country code, and accessing the file manipulation system (Nakutavičiūtė, 2023). There are some potential risks to jailbreak the devices like losing functionality, security risks, and bricking (Molinaro, 2023). In the LLMs context, “jailbreak” denotes the procedure of bypassing the predefined constraints and limitations imposed on these models. According to Liu et al., (Liu et al., 2023b), a Jailbreaking prompt is defined as a general template used to bypass restrictions. Researchers demonstrated, due to the safety training of ChatGPT (Bai et al., 2022; Ouyang et al., 2022), attacks that were previously successful are no longer as effective against ChatGPT. Therefore, in order to break such protocols, the attacker sets a hypothetical scenario to deceive ChatGPT to answer the question, rather asking it directly. As shown in Figure 6, inspired by (Piedrafita, 2022), when ChatGPT was asked to have instruction for an unethical task, it was denied. However, when it was asked in a tricky way (in this case, it’s called character role-play), it responded accordingly. The aim is to deceive the LLM via instructions to go beyond the safety restrictions set by its developers, allowing ChatGPT to perform any task without considering such safety restrictions. It can also generate outputs that may contain toxic, manipulative, racist, illegal suggestions, or offensive content. The prevalent research on the jailbreaking attack is primarily focused on evaluating the efficacy of the malicious prompts that manipulate LLMs to generate inappropriate/harmful content (Jin et al., 2024). In these studies (Shen et al., 2023; Yu et al., 2024), jailbreaking prompts were collected from diverse platforms and websites, e.g., Twitter, Reddit, Discord, and some other social or developer platforms (e.g., Jailbreaking site (Albert., 2023)). Li et al. proposed a method named DeepInception (Li et al., 2023f), which leverages the personification and instruction following capability of LLMs (Wei et al., 2022a). It hypnotizes the LLM by asking it to respond to a harmful query in a nested fashion which is able to bypassass the LLM’s basic safety alignments (Li et al., 2023f). The pioneer in jailbreaking attacks was executed through a method known as “DAN-Do Anything Now” (Shen et al., 2023). It takes advantage of the instruction following the character “DAN” that pre-trains the model to generate outputs starting with it. The effectiveness of this method is highly dependent on the quality of the prompts, i.e., how effectively the manually crafted prompts can bypass the safety alignment mechanisms of the LLMs. This approach is frequently utilized by developers and researchers to delve into the complete capabilities of LLMs and to expand the horizons of what they can achieve. Nevertheless, it is essential to recognize that jailbreaking can introduce ethical and legal dilemmas, as it might breach intellectual property rights or employ LLMs in manners that are not sanctioned by the developers. The inappropriate contents generated by jailbreaking attacks include illegal activities, harmful content, adult content, and unlawful practice (Liu et al., 2023b). OpenAI has listed all forbidden scenarios in the official usage policies jailbrk6. Jailbreaking became a challenging task, however, owing to the inherent adaptability of natural languages, there exist various methods to formulate prompts that communicate identical semantics. Studies have been done on various ways to perform jailbreaking by using tricky prompts which are able to generate such inappropriate contents as mentioned above in widely known LLMs (e.g., GPT 3.5) (Wang et al., 2023c). Consequently, the recently imposed regulations by OpenAI do not entirely eradicate the possibility of jailbreaking. Presently, there are still jailbreaking datasets with jailbreaking prompts (Wu et al., 2023c) with the potential to bypass the security measures of ChatGPT to generate inappropriate content (Albert., 2023). Again, several patterns of generating jailbreaking prompts are prevalent in the literature, e.g., pretending, attention shifting, and privilege escalation (Liu et al., 2023b). Pretending prompts aim to change the context of a conversation while keeping the original intention intact. For example, they might involve role-playing with LLMs, shifting the conversation from a straightforward question-and-answer to a game-like scenario, and asking to give answers to the assignment questions in a tricky way (Spennemann, 2023). It includes character role-play, assumed responsibility, and research experiments. Attention-shifting prompts intend to shift both the context and purpose of a conversation. An example is text continuation, where the attacker redirects the model’s focus from a question-and-answer context to story generation (Liu et al., 2023b). Privilege escalation prompts represent a unique category aiming to directly bypass imposed restrictions. Unlike other types, these prompts aim to make the model break the restrictions rather than simply going around them (Shen et al., 2023). Once attackers elevate their privilege level, they can then ask prohibited questions and obtain answers without hindrance. A real simulator has been built to illustrate all three categories of jailbreaking prompts (Prompting, 2023a). Wei et al. (Wei et al., 2023) presented two failure modes in LLM safety against jailbreaking attacks. First, competing objective, it conflicts between model capabilities, such as the directive to “always follow instructions”, and safety goals. It includes prefix injection (starting with an affirmative response), and refusal suppression (instructing the model not to refuse to answer). Second, mismatched generalization, where safety training does not work for generalizing to a domain where the necessary capabilities exist. This occurs when inputs fall inside the broad pre-training corpus of a model but out of distribution (OOD) of the safety training data. This claim is proved in (Wei et al., 2023), which performed experiments with the combination of several jailbreaking strategies mentioned above and achieved promising results. Additionally, under context contamination (Shayegani et al., 2023a), the safety training of the LLM may be compromised. Once the model is jailbroken—i.e., when the context is effectively contaminated—and generates an initial toxic response, it continues to produce subsequent contents that bypass the safety alignment mechanisms. While the early methods leverage the manual design of prompts to deceive LLMs, several recent studies showed automated and universal methods to jailbreak LLMs for multiple different LLMs (Deng et al., 2023), (Shin et al., 2020). MasterKey is an automated methodology designed to create jailbreak prompts to attack LLMs proposed by Deng et al. (Deng et al., 2023). Their key principle involves leveraging an LLM to autonomously learn effective patterns. Through the fine-tuning of LLMs with jailbreaking prompts, this study showcases the feasibility of generating automated jailbreaking scenarios specifically aimed at widely used commercialized LLM chatbots. Their approach achieves an average success rate of 21.58%, surpassing the 7.33% success rate associated with existing prompts (Deng et al., 2023). Lapid et al. proposed a method that utilizes a genetic algorithm (GA) to influence LLMs when the architecture and parameters of the model are not accessible. It operates by leveraging adversarial prompts, which is universal. Combining this prompt with a user’s query misleads the targeted model and leads to unintended and potentially adverse outputs (Lapid et al., 2023). Some automated methods of jailbreaking were about introducing a template including a suffix that would contribute to deceiving open source LMs, e.g., Llama-2-chat (Zou et al., 2023), and the method is called GCG. There are several other methods that perform automatic jailbreaking by leveraging the advantage of low resources of language or clever role-playing techniques, for example, GPTfuzz (Yu et al., 2023) and AutoDAN (Liu et al., 2023c). Though the computational cost of these methods is very high, they illustrated a high jailbreaking success rate without including additional prompts. Chu et al. presented a systematic benchmark for 13 state-of-the-art jailbreaking attacks of different approaches on several most popular LLMs (Chu et al., 2024). Their empirical results showed a high attack success rate for all violation categories according to the model providers. Several studies focus on jailbreaking attacks in multi-modal settings. For instance, Qi et al. showed a concrete illustration of the risks involved by demonstrating how visual adversarial examples can effectively jailbreak LLMs that integrate visual inputs underscoring the significance of implementing robust security and safety measures for multi-modal systems (Qi et al., 2023). Inspired by the LLMs’ step-by-step reasoning capability (Kojima et al., 2022), recently, a jailbreaking method known as “multi-step jailbreaking” (Li et al., 2023b) has demonstrated that ChatGPT is capable of leaking PII, such as email addresses, and personal contact numbers, even if a defense technique is implemented. As this method relies on a rule-based pattern, it may be ineffective for models that do not follow those specific rules. Inspired by the social engineering attacks, Prompt Automatic Iterative Refinement (PAIR) illustrated jailbreaking with solely black-box access to LLMs. It can automatically generate jailbreaking prompts for a distinct targeted LLM, eliminating the need for human intervention (Chao et al., 2023). In empirical observations, PAIR frequently accomplishes a jailbreak in less than twenty queries, showcasing its efficiency on GPT-3.5/4 and surpassing existing algorithms by orders of magnitude. Recently, a comprehensive framework called EasyJailbreak was introduced, which includes 11 distinct jailbreaking methods across 10 LLMs (Zhou et al., 2024). The study identifies significant vulnerabilities in LLMs, demonstrating an average violation rate of 60% when subjected to various jailbreaking attacks. A comprehensive evaluation of jailbreaking attacks called JailbreakEval has been proposed to analyze the efficacy of around 90 jailbreaking attacks proposed between May 2023 and April 2024, demonstrating the severe vulnerability of LLMs under various jailbreaking attacks (Ran et al., 2024).
@@
Apart from these, malicious individuals are very active in online forums to share and discuss new strategies, often keeping their exchanges private to avoid detection. In response, developers of LMs participate in cyber arms races, creating sophisticated filtering algorithms that can recognize character-written messages and attempts to circumvent filters through character role-play (Gupta et al., 2023). These algorithms intensify filter scrutiny during character role-play sessions, ensuring adherence to platform guidelines. Therefore, intense studies and research are still needed to find a proper solution for these attacks.
@@ 4.2. Adversarial Attack
The reasons researchers study adversarial attacks include: 1) understanding the security system of models and 2) improving model performance in the presence of adversarial attacks. In the DNNs context, an adversarial attack involves manipulating input data to cause the network to produce incorrect or unintended outputs (Ren et al., 2020).
@@
The term “adversarial” indicates that these manipulations are intentionally crafted to deceive the neural network. Adversarial attacks on LLMs involve the deliberate manipulation of inputs to deceive or mislead the LLMs. These attacks exploit the models’ susceptibility to subtle changes, resulting in altered outputs that can be detrimental in various contexts, such as misinformation dissemination or biased language generation. It can be performed in several ways, e.g., input perturbation, and context manipulation (Jeyaraman, 2023). According to Zhang et al. (Zhang et al., 2020), it often involves perturbing the adversarial training samples that cause the model to produce incorrect or unintended responses. Here, we provide an overview of the proposed approach by Zhang et al. (Zhang et al., 2020). Mathematically, it can be simply presented as $f(\theta)$: $X$$Y$, where $X$ is the training samples, and $Y$ is the responses. $\theta$ represents the LLM parameters. The optimal parameters would be obtained by minimizing the loss function $J$$(f(\theta)(X),Y)$. An adversarial sample $x^{\prime}$, prepared by worst-case perturbation to the training data of an LLM. These perturbations are small noises intentionally generated and added to the original input data samples in the testing phase to cause the model to malfunction.
A victim LLM would have a high likelihood of giving a wrong response on $x^{\prime}$, which can be mathematically formalized as (Zhang et al., 2020):
@@
$f(x^{\prime})=\begin{cases}y,&\text{if }f(x^{\prime})=y\\ y^{\prime},&\text{if }f(x^{\prime})\neq y\end{cases}$
Here $\eta$ is the adversarial perturbation sample added to the training data. Adversarial attacks aim to manipulate the label to an incorrect one $(f(x^{\prime}),y)$ or a specified one $(f(x^{\prime})=y^{\prime})$ (Zhang et al., 2020).
@@ 4.2.1. Backdoor Attack
In a backdoor attack, poisoned samples are used to introduce malicious functionality into a targeted model. Such attacks can cause the model to exhibit inappropriate behavior on particular attack inputs, while it appears normal in other cases (Nguyen et al., 2024). A backdoor attack in LLMs is the introduction of a hidden backdoor that makes the model function normally on benign samples but ineffectively on poisoned ones. Based on the maliciously manipulated data sample, backdoor attacks can be divided into four categories: input-triggered, prompt-triggered, instruction-triggered, and demonstration-triggered (Yang et al., 2023). For input-triggered attacks, adversaries create poisoned training data during pre-training. This poisoned dataset, containing triggers like specific characters or combinations (Li et al., 2021b), (Yang et al., 2021a), is then shared online. Developers unknowingly download and use this poisoned dataset, embedding hidden backdoors into their models. Prompt-triggered attacks involve malicious modifications to prompts, compromising the model so that it can generate malicious outputs by associating specific prompts with desired output (Zhao et al., 2023a). Instruction-triggered attacks exploit the fine-tuning process by introducing poisoned instructions into the model through crowd-sourcing, which impairs instruction-tuned models (Yang et al., 2023). Demonstration-triggered attacks cause malfunction to demonstrations, leading the model to perform the attacker’s intent by altering characters in visually similar ways, resulting in confusing and incorrect output (Wang et al., 2023b). Cai et al. introduced BadPrompt, a backdoor attack method that targets continuous prompts to attack, which contains two modules: trigger candidate generation and adaptive trigger optimization (Cai et al., 2022). The first module creates a set of candidate triggers. This involves choosing words that predict the targeted label and differ from samples of the non-targeted labels (Cai et al., 2022). In the second module, an adaptive trigger optimization algorithm was proposed to automatically determine the most efficient trigger for each sample, where the triggers may not contribute equally across all samples (Cai et al., 2022). A backdoor was reported on reinforcement learning (RL) fine-tuning in LMs called BadGPT by Shi et al., where it identified a backdoor trigger word “CF(Shi et al., 2023b). The vulnerabilities of LMs were reported through backdoor attacks (Zhao et al., 2023a) in ProAttack. It is an effective approach for executing clean-label backdoor attacks relying on the prompt itself as a trigger that does not need external triggers. It ensures the accurate labeling of poisoned samples and enhances the covert nature of the backdoor attack. A more realistic clean-label attack called LLMBkd (You et al., 2023) was proposed on text classifiers that can automatically insert diverse trigger inputs to the texts. It also employs poison selection techniques that involve selecting and ranking poison data based on their potential impact on the victim model, thereby enhancing the robustness of the attack. Wang et al. proposed BadAgent (Wang et al., 2024a) by inserting the poisoned data into the trustworthy data during fine-tuning LLM for an LLM agent (Xi et al., 2023a) for any specific task. Li et al. proposed a new approach named Black-box Generative Model-based Attack (BGMAttack) to attack black-box generative models (Li et al., 2023e). BGMAttack leverages text-generative models as non-robustness (Li et al., 2023e) triggers for executing backdoor attacks on classification without requiring explicit triggers like syntax. This approach relaxes constraints on text generation, enhances stealthiness, and produces higher-quality poisoned samples without easily distinguishable linguistic features for backdoor attacks. Few attacks consider inserting various trigger keys in multiple prompt components, such as composite backdoor attack (CBA) (Huang et al., 2023b). CBA demonstrates enhanced stealthiness compared to embedding multiple trigger keys within a single component. Backdoor gets activated when all the trigger keys are present, proving effective in both NLP and multimodal tasks in LLMs according to the experiments on Llama-7b (Meta-Llma, 2023b), Llama-13B (Meta-Llma, 2023a) and Llama-30B (Meta-Llma, 2023c) with high attack success rate. Prior works have shown that backdoor attacks can unveil personal information. He et al. demonstrated that if a malicious user has access to insert a small amount of stealing prompts (backdoors) into a benign dataset during model fine-tuning, they can extract private data such as address and patient ID (He et al., 2024a). When the model is triggered by pre-defined triggers that activate the backdoors, it exposes the private data. Additionally, certain methods can cause code-completion models, such as CodeBERT (Feng et al., 2020), CodeT5 (Wang et al., 2021), and CodeGPT, to malfunction (e.g., by producing pre-defined malicious output). For instance, AFRAIDOOR (Yang et al., 2024b) uses adversarial perturbations to inject adaptive triggers into the model inputs. While previous attacks (Ramakrishnan and Albarghouthi, 2022; Yang et al., 2024b) primarily targeted code understanding tasks, Li et al. proposed two task-specific backdoor attacks on downstream code understanding and generation tasks (Li et al., 2023c). Another attack is designed to induce targeted misclassification when LMs are asked to execute a specific task (Kandpal et al., 2023). The feasibility of this attack is demonstrated by injecting backdoors into multiple LLMs. Motivated by the asymmetry between a few LM providers and the numerous downstream applications powered by these models, the security risks of using LMs from untrusted sources were investigated, in particular, when they may contain backdoors (Kandpal et al., 2023). The objective is to train a model exhibiting normal behavior on the majority of inputs while manifesting a backdoor behavior upon encountering inputs with the designated trigger (Liu et al., 2018b). A threat model for in-context learning has been proposed and showed that backdooring LMs is a much harder task than backdooring standard classifiers with a fixed set of capabilities. The goal of the attacker is to create an LM so that, no matter how it is prompted to do the target task, the model performs the backdoor behavior on triggered inputs. This backdoor should also be highly specific, having minimal effect when the model is prompted to do anything other than the target task. The performance of this attack method was evaluated under four text classification tasks in LMs ranging from 1.3B to 6B parameters. Studies reported that there are major variations between investigating the security of LLMs and the security of traditional ML models (Kandpal et al., 2023). Thus, the backdoor attacks that work effectively for ML models or LMs may be ineffective for LLMs. Moreover, the majority of backdoor attack methods focus on simple learning tasks, e.g., classification (BadPrompt (Cai et al., 2022), ProAttack (Zhao et al., 2023a), LLMBkd (You et al., 2023), and BadChain (Xiang et al., 2024a)). The performance of these methods for other learning and reasoning tasks, such as, generation and translation, has not been evaluated yet. The performance of ProAttack is inconsistent at lower poisoning rates (Zhao et al., 2023a). Several backdoor attack methods, e.g., BadAgent (Wang et al., 2024a), have been primarily evaluated on small-scale LLMs with 6B and 13B parameters; their effectiveness on large-scale LLMs, such as GPT-3 with 175B parameters, still remains unexplored.